OpenStudy (anonymous):

what is the simplest form of this expression 4(-5y+4)+(-7)(-3y+1)

OpenStudy (owlcoffee):

Apply the distributive property and operate similar terms, that'd give you the answer.

OpenStudy (anonymous):

need more help then that

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@welshfella

Nnesha (nnesha):

distribute parentheses by outside terms (as owl suggested )

Nnesha (nnesha):

here is an example \[\huge\rm \color{red}{a}(b+c)=\color{red}{a} \times b + \color{red}{a} \times c =\color{reD}{a}b+\color{reD}{a}c\]

Nnesha (nnesha):